OpenStudy (anonymous):

the number 2.69x10^(-3) is equivalent to (1) 0.6269 (2) 0.269 (3) 269 (4) 0.00269

OpenStudy (whpalmer4):

Do you know what \[10^{-3} \]represents? \[10^3 = 10*10*10 = 1000\]\[10^{-3} = \frac{1}{10^3} = \frac{1}{10*10*10}=\frac{1}{1000} = 0.001\] So armed with that knowledge, what do you think 2.69 * 10^-3 is?

OpenStudy (anonymous):